OpenStudy (anonymous):

Carlos bought a pizza with an area of approximately 201in. What is the diameter of the pizza?

OpenStudy (campbell_st):

use the area of a circle formula to find the radius r, then double your answer for the diameter \[201 = \pi r^2\] \[r^2 = \frac{201}{\pi}\] \[r = \sqrt{\frac{201}{\pi}}\] evaluate and double the value to find the diameter.